OpenStudy (anonymous):

How do I prove this?????? 1 + tanx (tan(x/2)) = secx

OpenStudy (mertsj):

\[1+\frac{\sin x}{\cos x}(\frac{1-\cos x}{\sin x})=1+\frac{1-\cos x}{\cos x}=\]

OpenStudy (mertsj):

\[1+\frac{1}{\cos x}-\frac{\cos x}{\cos x}=1+\sec x-1=\sec x\]

OpenStudy (anonymous):

I'm not understanding what you just did Could you explain?

OpenStudy (mertsj):

I replaced tanx with sin/cos I replaces tan(x/2) with (1-cosx)/sinx I cancelled sinx from numerator and denominator I split the fraction into two fractions I replaced 1/cosx with secs. I replaced cosx/cosx with 1 I subtracted 1-1

OpenStudy (anonymous):

Awesome, would this be the standard method for proving problems or is it just for this problem?

OpenStudy (mertsj):

It would be standard form for proving identities.

OpenStudy (anonymous):

So just to be clear I can use the steps you just provided for proving all identities?

OpenStudy (mertsj):

No. Only the ideas transfers to another problem. The ideas of making the proper substitutions and simplifying and of working on one side only. Every problem is unique.

OpenStudy (anonymous):

Ok